the osmotic pressure of blood is 7.65atm at 37 degre celsius how much glucose should be used per liter for an intravinuous injection that is to have the same osmotic pressure as blood and R=.00821
1
Expert's answer
2014-05-05T04:36:23-0400
π = cRT
T = 273 + 37 = 310
c = π/RT = 7.65/(0.00821*310) = 3.01 M